a painter leans a 34 ft ladder on a building. the base of the ladder is 16 ft from the building. to the nearest foot, how high on the building does the ladder reach?
please help me.

Answer by uday1435(57)

Here the ladder forms the hypotenuse of a right triangle. Let the height be x
By Pythagorean theorem 34^2 = 16^2 + x^2
That is 1156 = 256 + x^2
So x^2 = 1165 – 256 = 900
x = √900 = 30 ft
Ans: The ladder reaches 30 ft high



If you still require further clarifications please write to udayakumar.t.r@gmail.com